OpenStudy (anonymous):

Given the function f(x) = 4x-2/3, which of the below expressions is correct?

OpenStudy (anonymous):

\[f^-1(x)=\frac{ 2-4x }{ 3 }\] \[f^-1(x)=\frac{ 3x -2}{ 4 }\] \[f^-1(x)=\frac{ -4x-2}{ 3 }\] \[f^-1(x)=\frac{ 3x+2}{ 4 }\]

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

Assuming you meant (4x-2)/3: To do inverse functions, swap x and y: \[\frac{ 4y-2 }{ 3 }=x\] Solve for y: \[3x =4y-2\]\[3x+2=4y\]\[y=\frac{ 3x+2 }{ 4}\] Option 4 is correct.

OpenStudy (anonymous):

Thank you!